OpenStudy (anonymous):

Leo paid $2400 in interest on an amount borrowed for 10 years at a 4% annual simple interest rate. How much did Leo borrow?

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

o thanks i think it is 6000

OpenStudy (anonymous):

???

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (dan815):

simple interest so its agreed u just pay 4% of the 2400 every year so 0.04 * 2400 = interest for 1 year you do that for 10 years so 10*0.04*2400

OpenStudy (anonymous):

ok so =96

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

RELLY WHY DOES PEOPLE DO THIS TO ME I HATE IT WHEN YOU DOIT TO ME ALL I NEED IS THIS ONE QUESTION!!!!!?

OpenStudy (anonymous):

@dan815

TheSmartOne (thesmartone):

How did you get 96?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks for helping me

TheSmartOne (thesmartone):

\(\sf I = PRT\) Where I is interest P is the amount borrowed R is the interest rate T is the time.

OpenStudy (anonymous):

ok so is it like pemdas?

TheSmartOne (thesmartone):

So they said that the interest rate is 4% or .04 They also said that the interest is 2400 so I = 2400 and T = 10 years

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

Well basically all we have to do is use the formula and just solve for it :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so we multiply

TheSmartOne (thesmartone):

So plugging in those numbers: \(\sf I = P \times R \times T\) Plug in the numbers: \(\sf 2400 = P \times .04 \times 10\) Solve for P

OpenStudy (anonymous):

so it is p=4

TheSmartOne (thesmartone):

Actually no.

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

Solve for P \(\sf 2400 = P \times .04 \times 10\) So we need to divide by .04 and 10 on both sides so: \(\sf\Large\frac{2400}{.04 \times 10} = \frac{P \times .04 \times 10}{.04 \times 10}\) And if we simplify that we get \(\sf\Large P = \frac{2400}{.04 \times 10}\)

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

And btw, @dan815 was wrong lol :P He thought that the P = 2400, when it actually is I = 2400

OpenStudy (anonymous):

ok so it is 2400

TheSmartOne (thesmartone):

No, do the multiplication in the denominator and then divide. What is \(\sf .04 \times 10=?\)

OpenStudy (anonymous):

it is 0.4

TheSmartOne (thesmartone):

Correct. So then what is \(\sf\Large\frac{2400}{.4}=?\)

OpenStudy (anonymous):

it is 600

OpenStudy (dan815):

oh oops :) sry!!

OpenStudy (dan815):

i did not read the question

TheSmartOne (thesmartone):

No, use a calculator. If you don't have one, then remember that \(\sf\Large .4 = \frac{4}{10}\) So then that means \(\sf\Large\frac{2400}{.4}=\frac{2400}{\frac{4}{10}} = 2400 \div \frac{4}{10}=2400 \times \frac{10}{4}= \frac{2400 \times 10}{4}\)

OpenStudy (anonymous):

yea thanks @dan815

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

So what did you get? 600 is wrong.

OpenStudy (anonymous):

ok it is 6000

OpenStudy (anonymous):

right?

TheSmartOne (thesmartone):

Correct! :)

OpenStudy (anonymous):

ok so that is the answer

TheSmartOne (thesmartone):

It is (:

OpenStudy (anonymous):

ok thanks for stepingin and helping me other than dan815 thanks for saving the day

TheSmartOne (thesmartone):

Anytime :D

OpenStudy (anonymous):

have a nice day

TheSmartOne (thesmartone):

I don't respond to tags really fast because I get to many of them. But if you need my help fast, tag me and also message me. Have a great day (:

OpenStudy (anonymous):

ok bye